OpenStudy (anonymous):

Find dy/dx in terms of t for the following curve: (c) x=2cost, y=3sint

OpenStudy (dumbcow):

\[\frac{dy}{dx} = \frac{dy}{dt}*\frac{dt}{dx}\]

OpenStudy (anonymous):

True, but I'm not good with cos and sin How do I work with it? Would x' = -2sint y'= 3cost ??

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

so, it would be\[ 3cost \over -2sint? \]

OpenStudy (dumbcow):

yep or -3/2 cot(t)

OpenStudy (anonymous):

Is cos/sin always cot?

OpenStudy (dumbcow):

yes and sin/cos is always tan

OpenStudy (anonymous):

Yes, I knew that one :) Do you think you could help me with the equation I just put up?

OpenStudy (dumbcow):

yeah i'll take a look